OpenStudy (pphalke):

If a 10 foot ladder is leaned against the top of a 9 foot high wall, how far will the base of the ladder be from the bottom of the wall? round your answer to the nearest 10th of a foot

OpenStudy (daniellelovee):



OpenStudy (daniellelovee):

do you know pythagorean theory? a^2+b^2=c^2

OpenStudy (daniellelovee):

right now we have c=10 a=9 and b=?

OpenStudy (daniellelovee):

can you solve from there?

OpenStudy (pphalke):

i think so i am not sure how

OpenStudy (daniellelovee):

ok then try it and I will revice if you do something incorrectly :)

OpenStudy (pphalke):

i got 181 by doing 9^2+10^2

OpenStudy (daniellelovee):

that is wrong you have to do 9^2+b^2=10^2 so you would subtract not add

OpenStudy (pphalke):

how do i do that?

OpenStudy (daniellelovee):

you have to follow Pythagorean theory

OpenStudy (daniellelovee):

you have to find b

OpenStudy (daniellelovee):

do you understand?